8. 2 sinº x-3 sin x-1=0 11. 4 sin’ x + 7 sin x = 6 9. tanx-1=0 12. ( tan x-)(tan x-3) = 0
Answer #1

③ 2 sins-3sina-120 Let sinaay => 2y2 -38-130 Y = x3 = 99 +8 = 32517 4 4 Now sina = 3=SIT .. 3+517 yo 4 Not valid + sind=3-J17(12 (+ang - ) ( x - 3): tana and tanx=3 a) tanda tan 71.565 4 staractament 3) 2= 71.565 7) a= tant = 26.565